OpenStudy (anonymous):

How do I factor 5 sq root 24 times 3 sq root 10

OpenStudy (anonymous):

\[5\sqrt{24} \times 3\sqrt{10}\]

OpenStudy (anonymous):

If you are looking for the answer to (5sqrt24) x (3sqrt10) the answer is 54,000

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

i got this by multiplying

jimthompson5910 (jim_thompson5910):

It might help to simplify first before you multiply. So let's simplify \(\large \sqrt{24}\) first. \[\large \sqrt{24} = \sqrt{4*6}\] \[\large \sqrt{24} = \sqrt{4}*\sqrt{6}\] \[\large \sqrt{24} = 2\sqrt{6}\] Notice how I factored 24 into 4*6 and 4 is a perfect square.

jimthompson5910 (jim_thompson5910):

Since 10 doesn't have any factors which are perfect squares (other than 1), this means that \(\large \sqrt{10}\) cannot be simplified

jimthompson5910 (jim_thompson5910):

So that means we'll have this \[\large 5\sqrt{24} \times 3\sqrt{10}\] \[\large 5*2\sqrt{6} \times 3\sqrt{10}\] \[\large 10\sqrt{6} \times 3\sqrt{10}\] \[\large (10*3)*(\sqrt{6}*\sqrt{10})\] \[\large 30\sqrt{6*10}\] I'll let you finish up.

OpenStudy (anonymous):

Now if I have \[2\sqrt{6}\] can i multiply it with \[3\sqrt{10}\]?

jimthompson5910 (jim_thompson5910):

yes you can

jimthompson5910 (jim_thompson5910):

you do so by multiplying the outer coefficients together

jimthompson5910 (jim_thompson5910):

then combining the roots using the idea that \(\large \sqrt{x}*\sqrt{y} = \sqrt{x*y}\)

OpenStudy (anonymous):

\[6\sqrt{60}\]

jimthompson5910 (jim_thompson5910):

Don't forget about the 5 as well

OpenStudy (anonymous):

60= 2,30= 2,15 = 3,5 2 \[\sqrt{60}=2\sqrt{15}\]

OpenStudy (anonymous):

what five?

OpenStudy (anonymous):

wait so the answer is 12 sq root 15?

jimthompson5910 (jim_thompson5910):

up at the top you have "5 sq root 24 times 3 sq root 10" so the first number in the expression

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

so it's 30 sq root 60?

jimthompson5910 (jim_thompson5910):

now simplify \(\large \sqrt{60}\)

OpenStudy (anonymous):

\[30\sqrt{60}=30(2\sqrt{15})=60\sqrt{15}\]

jimthompson5910 (jim_thompson5910):

you are correct, the final answer is \(\large 60\sqrt{15}\)

OpenStudy (anonymous):

could you help me with one more similar problem really quick?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

\[\sqrt{3}\times \sqrt{27}\]

jimthompson5910 (jim_thompson5910):

I would simplify \(\large \sqrt{27}\) first

jimthompson5910 (jim_thompson5910):

or I guess you could just combine the roots and multiply

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

so i would just make it \[\sqrt{3}\times3\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

then you combine the roots and multiply

OpenStudy (anonymous):

which is just 3 * 3 =9

jimthompson5910 (jim_thompson5910):

then take the square root of that to get 3

OpenStudy (anonymous):

oh snap I forgot about that okay that makes sense

jimthompson5910 (jim_thompson5910):

that's ok, i have brain farts all the time too

OpenStudy (anonymous):

could you help me with one last question? i swear thats it after this

jimthompson5910 (jim_thompson5910):

sure I'd love to

OpenStudy (anonymous):

\[\frac{ 3 + \sqrt {7}}{ 2 - \sqrt 10 }\]

OpenStudy (anonymous):

i know i have to multiply the top and bottom by 2 + sqrt 10

jimthompson5910 (jim_thompson5910):

good start, what does that give you once you do?

OpenStudy (anonymous):

\[\frac{ 6+ 3\sqrt10 +2\sqrt7+\sqrt70 }{ 4 - \sqrt10 + \sqrt10 - \sqrt100}\]

OpenStudy (anonymous):

And then i'd combine all the like terms

jimthompson5910 (jim_thompson5910):

so far, so good

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

to get \[\frac{ 6 + 3\sqrt7+2\sqrt7+\sqrt70 }{ -6 }\]

OpenStudy (anonymous):

and that's where I'm stuck

jimthompson5910 (jim_thompson5910):

what are the like terms in this case?

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (anonymous):

sqrt 7?

jimthompson5910 (jim_thompson5910):

correct, if you replace \(\large \sqrt{7}\) with z (or any other variable really), you go from \[\large 3\sqrt7+2\sqrt7\] to \[\large 3z + 2z\] and that looks much clearer in terms of spotting the like terms (and combining them)

OpenStudy (anonymous):

so would i multiply those two and get (3 * 2) (sqrt7 * sqrt7) = 6*7=42?

jimthompson5910 (jim_thompson5910):

nope, look at my last message

OpenStudy (anonymous):

wait no i wrote that wrong it's supposed to be \[\frac{ 6+3\sqrt10+2\sqrt7+\sqrt70 }{ -6 }\]

jimthompson5910 (jim_thompson5910):

oh right, not sure where the root 10 went...

jimthompson5910 (jim_thompson5910):

at this point, there's not much you can do

OpenStudy (anonymous):

I just wrote it down wrong. This is where I have trouble simplyfing it.

jimthompson5910 (jim_thompson5910):

because there are no like terms in the numerator

OpenStudy (anonymous):

what bout the 6 and -6?

OpenStudy (anonymous):

*about

jimthompson5910 (jim_thompson5910):

I guess you could go from \[\frac{ 6+3\sqrt{10}+2\sqrt{7}+\sqrt{70} }{ -6 }\] to \[\frac{ -6-3\sqrt{10}-2\sqrt{7}-\sqrt{70} }{ 6 }\]

jimthompson5910 (jim_thompson5910):

but that's about all you can do really

OpenStudy (anonymous):

i thought you could get rid of the 6 and 6

jimthompson5910 (jim_thompson5910):

if you do that, you would have to break up the fraction

OpenStudy (anonymous):

oh whoops

OpenStudy (anonymous):

okay that's it haha thank you so much for all the help I really appreciate it

jimthompson5910 (jim_thompson5910):

You're welcome, I'm glad I could be of help.

OpenStudy (anonymous):

Have a nice night or day wherever you are

jimthompson5910 (jim_thompson5910):

thanks, you too